OpenStudy (anonymous):

Graph each equation on your graphing calculator and identify the vertex of each function: 1. y=4(x-2)^2 2. f(x)=(x+1)^2 +2 3. . f(x) = x^2 - 4x + 5 4. . f(x) = -2x^2 + 4x - 3 5. y = x^2 + 5x - 14 6. A ball is dropped from the top of a building. The distance in meters above the ground y of the ball after t seconds can be modeled by the equation y = -9.8t^2 + 100. a. What is the y-intercept of the equation? b. Describe the meaning of the y-intercept of the graph of the equation. Factor each expression: 7. x^2 + 3x - 54 8. x^2 + 10x + 24 9. 25x^2 - 36 10. 3x^2 - 9x

OpenStudy (anonymous):

11. x^2 - 15x + 56 12. 4x^2 + 28x + 49 Solve each equation by factoring: 13. 2x^2 + 4x = 0 14. . 4x^2 - 100 = 0 15. x^2 – 2^x + 1 = 0 16. x^2 - 9 = 0 17. x^2 - 6x - 16= 0 18. 5x^2 + 23x - 10 = 0

OpenStudy (anonymous):

pleasse help!!